1. Let v be the vector (3 over p), such that |v| = 17. Calculate all possible values of p.

Answer by stanbon(75887)

Let v be the vector (3 over p), such that |v| = 17. Calculate all possible values of p.
----
|3/p| = 17
3/p = 17 or 3/p = -17
p = 17/3 or p = -17/3
-------
Cheers,
Stan H.
----------